# Star Attention: Efficient LLM Inference over Long Sequences

### v1

@@ Abstract
Inference with Transformer-based Large Language Models (LLMs) on long sequences is both costly and slow due to the quadratic complexity of the self-attention mechanism. We introduce Star Attention, a two-phase block-sparse approximation that improves computational efficiency by sharding attention across multiple hosts while minimizing communication overhead. In the first phase, the context is processed using blockwise-local attention across hosts, in parallel. In the second phase, query and response tokens attend to all prior cached tokens through sequence-global attention. Star Attention integrates seamlessly with most Transformer-based LLMs trained with global attention, reducing memory requirements and inference time by up to 11x while preserving 95-100% of accuracy.
@@ 1 Introduction
Recent Large Language Models (LLMs) can support contexts up to millions of tokens in length (Gemini-Team, 2024; Anthropic, 2024; Meta-AI, 2024), unlocking applications such as repository-level code analysis, multi-document summarization, and large corpus retrieval. However, processing such long sequences with LLMs requires substantial computational and memory resources due to the quadratic complexity of the self-attention mechanism.
To address these challenges, various techniques have been proposed to reduce memory usage and increase inference speed. For example, Flash Attention introduces an efficient GPU block-wise implementation of the global attention, achieving significant reductions in memory overhead and runtime (Dao et al., 2022; Dao, 2024). Ring Attention further extends this idea by distributing the computation of self-attention and feed-forward modules across multiple devices, cleverly overlapping communication with shard-local attention computations to enhance scalability (Liu et al., 2024a; Beltagy et al., 2020). More broadly, distributed strategies such as tensor, pipeline, sequence, and data parallelism have been proposed to divide compute effectively across multiple machines (Shoeybi et al., 2019; Huang et al., 2019; Li et al., 2023; Meta-AI, 2021).
Several prior works have shown that the attention matrix can be approximated with sparse attention mechanisms reducing the algorithmic complexity from quadratic to linear or log-linear. Child et al. (2019) significantly reduces the complexity of attention by leveraging sparse factorizations and (Choromanski et al., 2021) approximates attention using kernel-based methods. (Beltagy et al., 2020) employs sliding window attention and global tokens for efficient long-sequence processing while Xiao et al. (2024) adapts it for real-time long-sequence generation utilizing attention sinks. Complementing these approaches, memory-efficient techniques have also emerged. Key-value (KV) cache compression (Dai et al., 2019; Ge et al., 2024; Munkhdalai et al., 2024; Sun et al., 2024; Liu et al., 2024b) and low-rank approximations (Srebro & Jaakkola, 2003) trade precision for reduced memory usage.
We introduce Star Attention, a novel algorithm for efficient LLM long-context inference  [^1] Code is available at: https://github.com/NVIDIA/Star-Attention. This method is based on the observation that LLM inference usually has two stages: (1) prompt encoding, where the model processes input and stores KV vectors in the cache and (2) token generation, where model attends to the KV cache and autoregressively generates new tokens while updating the cache with the new KV vectors. In many long-context tasks, the input consists of a long context followed by a short query and a short answer. The information needed for answering the query is often localized within small parts of the context, meaning context tokens need only attend to nearby tokens, while query tokens need to attend to all prior tokens. Based on this observation, Star Attention utilizes a two-phase approach shown in Figure 1:
@@
    Context Encoding: The context is divided into contiguous blocks and distributed across “context” hosts, with each host also receiving a copy of the first block (an “anchor block”). Hosts compute self-attention only for their assigned blocks, without communicating with each other, reducing attention complexity from quadratic to linear with respect to context length. This distributed processing is similar to Ring Attention (Liu et al., 2024a) but without the “ring” communication during context encoding (Figure  1(a)).
@@
Star Attention enables the context length to scale linearly with the number of hosts by distributing the context processing across multiple hosts. Star Attention is compatible with most Transformer-based LLMs trained with global attention, operating seamlessly out-of-the-box without additional model fine-tuning. We evaluate Star Attention for Llama3.1-8B and Llama3.1-70B (Meta-AI, 2024) on several long-context benchmarks. Star Attention achieves up to 11 times faster inference while maintaining 95-100% of the baseline accuracy. Furthermore, Star Attention can be combined with other LLM optimization methods like Flash Attention or KV cache compression, allowing for additional speedup enhancements during inference.
@@
Figure 1: Star Attention inference flow. All devices in the system are grouped into hosts where one of the hosts is labeled as the “query” host. The input sequence is processed in two phases. Phase 1 - context encoding. The context portion of the input is partitioned into smaller blocks and distributed across hosts. All blocks, except the first, are prefixed with the initial block, called the “anchor” block. Each host processes its assigned block and stores the non-anchor portion of the KV cache. Phase 2 - query encoding and token generation. The input query is broadcast to all the hosts, where in each host, it first attends to the local KV cache computed during phase one. Then the “query” host computes global attention by aggregating the softmax normalization statistics from all the hosts. This process is repeated for each generated token.
@@
Figure 2: Block sparsity pattern for a sequence partitioned into 5 context blocks $c_{i}$ and a query block $q$. Each context block attends only to itself and the “anchor block” whereas the query attends to the entire input.
2 Star Attention Algorithm
2.1 Phase 1: Context Encoding
$c^{\prime}=[c_{1},(c_{1}\>c_{2}),(c_{1}\>c_{3}),\ldots,(c_{1}\>c_{n})]$
where each augmented block $c^{\prime}_{i}$ contains $2b$ tokens: $b$ tokens from the anchor block $c_{1}$ followed by $b$ tokens from the current block $c_{i}$ (Figure  1). The positional indices of $c_{1}$ are preserved, ensuring that its tokens retain their original position indices $[0,1,\ldots,b-1]$. The augmented blocks are distributed across compute hosts, where each host computes attention over the $2b$ tokens from its assigned block $c_{i}^{\prime}$ and generates the corresponding key-value (KV) vectors. While KVs for the anchor block $c_{1}$ are discarded, the KVs for the current block $c_{i}$ are retained in the cache.
@@ 2.1 Phase 1: Context Encoding
Figure 3: Attention distribution along the sequence length for context encoded with different strategies in phase 1 of Star Attention. (a) Global attention shows a spike at the start, corresponding to the attention sink. (b) Star Attention without anchor blocks shows several attention sinks present at the beginning of each block. (c) Star Attention with anchor blocks shifts sinks to anchor tokens, resulting in an attention distribution approximating global attention. In the plot, the input sequence (4K tokens) is divided into 512-token chunks.
We observe that, without anchor blocks—i.e., applying blockwise attention only to the original context $c$—the model fails to generate correct outputs. We conjecture this failure is due to the incorrect approximation to the attention patterns observed during phase 2 (Figure 3(b)), where multiple attention spikes, known as attention sinks (Xiao et al., 2024), are distributed across the sequence. These spikes occur because each block is processed independently, creating an attention sink at the start of each block. As a result, the model struggles to effectively focus on relevant parts of the context. To address this issue, we prefix the blocks with the anchor block $c_{1}$, shifting the attention sinks to the anchor tokens. By discarding the KVs of the anchor tokens the intermediate attention sinks are removed ensuring the attention distribution of block-local attention (Figure 3(c)) closely approximates global attention (Figure 3(a)) while maintaining the computational efficiency of blockwise processing.
@@ 2.2 Phase 2: Query Encoding and Token Generation
This method ensures that the global attention scores are normalized correctly across all hosts. It requires the communication of only a single scalar $s_{h}$ (the local sum of exponents) and a vector $A_{h}$ (the local attention) per token. In practice, the log-sum-exp method from online softmax (Milakov & Gimelshein, 2018) can be used to maintain the numerical stability during global attention aggregation. This distributed approach enables efficient computation by minimal data transfers between hosts.
@@
16 4 86.12 +2.47% 1.1x
32 8 82.52 +1.54% 1.2x
64 16 79.05 +1.28% 1.8x
Llama-3-8B-Instruct, 1048K Gradient.ai (2024) 128 32 77.39 +1.23% 2.7x
16 4 95.09 -2.85% 1.7x
32 8 94.61 -2.70% 2.0x
Llama-3.1-70B-Instruct, 128K Meta-AI (2024) 64 16 88.54 -1.63% 4.7x
Table 1: Star Attention vs Ring Attention (baseline) accuracy and relative inference speed-up. The $\Delta$ for Star Attention shows the relative accuracy improvement (+) or degradation (-). We set block size to one-quarter of the sequence length. Star Attention achieves significant speedup over Ring Attention while maintaining the accuracy. For larger models, the speedup of Star Attention is even more pronounced.
3 Experiments
We evaluate Star Attention on several Llama-based models with sequence lengths ranging from 16K to 1M tokens on RULER (Hsieh et al., 2024) and BABILong (Kuratov et al., 2024) benchmarks. We begin by comparing accuracy and the speed achieved by Star Attention versus baseline - Ring attention. Further, we investigate the impact of varying block sizes on accuracy, illustrating the trade-off between accuracy and speedup. Finally, we conduct a detailed analysis of challenging and favorable cases for Star Attention by examining distinct RULER task categories.
@@ 3.1 Setup
Models. We benchmark the base and instruct variants of the Llama-3.1 8B model which support context lengths up to 128K tokens (Meta-AI, 2024). In addition, we evaluate two Gradient.ai models that extend Llama-3-8B’s context up to 1M tokens Gradient.ai (2024). To access the scalability of our method, we also evaluate the Llama-3.1-70B-Instruct model. We observe that large LMs achieve even greater speedups with Star Attention on long context tasks.
Baseline. We compare Star Attention with Ring Attention (Liu et al., 2024a). Ring Attention computes global block-wise attention by having each host communicate its respective KV cache in a ring pattern across all the hosts . More details regarding our baseline selection in Appendix C.1.
Configuration. We implement Star Attention using HuggingFace Transformers library (Wolf et al., 2020). All experiments are done on A100 GPUs with bfloat16 precision. More details on the experiment configuration are in Appendix C.
Evaluation Benchmarks. We use the RULER benchmark for evaluation. It consists of 13 tasks categorized into 4 domains: Needle-in-a-Haystack (Retrieval), Multi-Hop Tracing, Aggregation, and Question Answering (Hsieh et al., 2024). Additionally, we report results on the BABILong benchmark, which encompass tasks where multiple supporting facts encoded in the context are required to generate accurate answers (Kuratov et al., 2024). Further details on the benchmarks and specific tasks can be found in Appendix B.
Figure 4: Accuracy (%) of Star Attention on RULER and BABILong evaluated on sequence lengths of 16K, 32K, 64K, and 128K. In all experiments, the block size and anchor block size are set to one-quarter of the total sequence length. Results using the Llama-3-8B-Instruct-262k, Llama-3.1-8B-Instruct and Llama-3.1-8B-Base models demonstrate that Star Attention retains 95-100% of the accuracy of global attention, and in some cases, even outperform it.
@@ 3.2 Results
Table 1 provides relative speedup and accuracy achieved by Star Attention versus Global (Ring) Attention from 16K to 128K tokens on the RULER benchmark. In each setting, the context block size and anchor block size are set to one-quarter of the total sequence length. Star Attention achieves similar accuracy to full global attention, with relative accuracy degradation limited to 0-3% while also giving upto 5x inference speedup. This demonstrates that Star Attention effectively preserves the model’s ability to retrieve relevant information, even with a significantly reduced context window. In case of larger models, such as Llama-3.1-70B Instruct, we find that these models achieves even greater speedups at any given sequence length while maintaining similar levels of accuracy degradation. We discuss Star Attention’s strengths and limitations based on RULER subtasks in Section 3.4. Full RULER scores for all models can be found in Appendix E.
Extending this analysis to other benchmarks and models, we evaluate Star Attention on the BABILong benchmark as well using Llama-3.1-8B-Instruct, Llama-3.1-8B-Base, and the gradientai-Llama-3-8B-Instruct-262K model. We have a similar observation here that Star Attention achieves similar accuracy to full global attention, with accuracy degradation limited to 0-3% across all tasks up to 128K tokens, as shown in Figure 4.
However, we observe several anomalies with the Llama-3.1 8B base model on the BABILong benchmark. There is a significant improvement at the 16K sequence length, but a severe drop at 128K. These fluctuations likely stem from the benchmark’s format-specific requirements for generating answers, which pose challenges for base models since they are not optimized for instruction-following tasks. As a result, scores for base models, especially at longer sequence lengths, may be less reliable. More details provided in Appendix D.

### v2

@@ Abstract
Inference with Transformer-based Large Language Models (LLMs) on long sequences is both costly and slow due to the quadratic complexity of the self-attention mechanism. We introduce Star Attention, a two-phase block-sparse approximation that improves computational efficiency by sharding attention across multiple hosts while minimizing communication overhead. In the first phase, the context is processed using blockwise-local attention across hosts, in parallel. In the second phase, query and response tokens attend to all prior cached tokens through sequence-global attention. Star Attention integrates seamlessly with most Transformer-based LLMs trained with global attention, reducing memory requirements and inference time by up to 11x while preserving 97-100% of accuracy.
@@ 1 Introduction
Recent Large Language Models (LLMs) can support contexts up to millions of tokens in length (Gemini-Team, 2024; Anthropic, 2024; Meta-AI, 2024; Qwen, 2025), unlocking applications such as repository-level code analysis, multi-document summarization, and large corpus retrieval. However, processing such long sequences with LLMs requires substantial computational and memory resources due to the quadratic complexity of the self-attention mechanism.
The importance of long-context capabilities has driven substantial research into addressing the computational challenges of self-attention. Some approaches focus on reducing the need to fully materialize the attention matrix (Milakov & Gimelshein, 2018), leading to blockwise processing techniques (Dao et al., 2022; Dao, 2024; Liu & Abbeel, 2024) and further optimization through distributed computation across multiple compute units (Liu et al., 2024a). While these methods improve training efficiency, autoregressive decoding during inference still requires the model to attend to every previous token, resulting in high computational costs for long-context sequences. Other approaches attempt to optimize inference by segmenting long inputs into chunks, encoding them separately, and attending to these encoded chunks using the user query (Beltagy et al., 2020; Russak et al., 2024; Liao et al., 2024). However, these methods often require fine-tuning the model or introducing additional components that necessitate further training, limiting their out-of-the-box applicability.
We introduce Star Attention [^1] Code is available at: https://github.com/NVIDIA/Star-Attention, a novel algorithm for efficient LLM long-context inference. This method is based on the observation that LLM inference usually has two stages: (1) prompt encoding, where the model processes input and stores KV vectors in the cache and (2) token generation, where model attends to the KV cache and autoregressively generates new tokens while updating the cache with the new KV vectors. In many long-context tasks, the input consists of a long context followed by a short query and a short answer. The information needed for answering the query is often localized within small parts of the context, meaning context tokens need only attend to nearby tokens, while query tokens need to attend to all prior tokens. Based on this observation, Star Attention utilizes a two-phase approach shown in Figure 1:
@@
    Context Encoding: The context is divided into contiguous blocks and distributed across “context” hosts, with each host also receiving a copy of the first block (an “anchor block”). Hosts compute self-attention only for their assigned blocks, without communicating with each other, reducing attention complexity from quadratic to linear with respect to context length. This distributed processing is similar to Ring Attention (Liu et al., 2024a) but without the “ring” communication during context encoding (Figure 1(a)).
@@
Star Attention enables the context length to scale linearly with the number of hosts by distributing the context processing across multiple hosts. Star Attention is compatible with most Transformer-based LLMs trained with global attention, operating seamlessly out-of-the-box without additional model fine-tuning. Furthermore, we combine Star Attention with Flash Attention, allowing for additional speedup enhancements. We evaluate Star Attention for Llama3.1-8B and Llama3.1-70B (Meta-AI, 2024) on several long-context benchmarks. Star Attention achieves up to 11 times faster inference while maintaining 97-100% of the baseline accuracy.
@@
Figure 1: Star Attention inference flow across two phases. (a) Context Encoding: The input context is partitioned into blocks and distributed across hosts, where each block (except the first) is prefixed with the anchor block ($c_{1}$). Each host processes its assigned block and stores the non-anchor portion of the KV cache. (b) Query Encoding and Token Generation: The query is broadcast to all hosts, which compute local attention using cached KVs. A designated “query” host then aggregates softmax normalization statistics to compute global attention and generates the next token.
2 Star Attention Algorithm
2.1 Phase 1: Context Encoding
$c^{\prime}=[c_{1},(c_{1}\>c_{2}),(c_{1}\>c_{3}),\ldots,(c_{1}\>c_{n})]$
where each augmented block $c^{\prime}_{i}$ contains $2b$ tokens: $b$ tokens from the anchor block $c_{1}$ followed by $b$ tokens from the current block $c_{i}$ (Figure  2.1). The positional indices of $c_{1}$ are preserved, ensuring that its tokens retain their original position indices $[0,1,\ldots,b-1]$. The augmented blocks are distributed across compute hosts, where each host computes attention over the $2b$ tokens from its assigned block $c_{i}^{\prime}$ and generates the corresponding key-value (KV) vectors. While KVs for the anchor block $c_{1}$ are discarded, the KVs for the current block $c_{i}$ are retained in the cache.
@@ 2.1 Phase 1: Context Encoding
Figure 2: Block sparsity pattern in Star Attention for a sequence partitioned into 5 context blocks $c_{i}$ and a query block $q$. Each context block attends only to itself and the “anchor block” whereas the query attends to the entire input.
@@
Figure 3: Attention distribution across the sequence during context encoding under different strategies in Phase 1. (a) Global attention exhibits a single attention sink at the sequence start. (b) Without anchor blocks, blockwise context encoding creates multiple attention sinks at the start of each block. (c) With anchor blocks, attention sinks shift to anchor tokens, yielding a distribution that closely approximates global attention. The sequence is 4K tokens long and partitioned into 512-token chunks.
We observe that, without anchor blocks—i.e., applying blockwise attention only to the original context $c$—the model fails to generate correct outputs. We conjecture this failure is due to the incorrect approximation to the attention patterns observed during phase 2 (Figure 3(b)), where multiple attention spikes, known as attention sinks (Xiao et al., 2024b), are distributed across the sequence. These spikes occur because each block is processed independently, creating an attention sink at the start of each block. As a result, the model struggles to effectively focus on relevant parts of the context. To address this issue, we prefix the blocks with the anchor block $c_{1}$, shifting the attention sinks to the anchor tokens. By discarding the KVs of the anchor tokens the intermediate attention sinks are removed ensuring the attention distribution of block-local attention (Figure 3(c)) closely approximates global attention (Figure 3(a)) while maintaining the computational efficiency of blockwise processing.
@@ 2.2 Phase 2: Query Encoding and Token Generation
This method ensures that the global attention scores are normalized correctly across all hosts. It requires the communication of only a single scalar $s_{h}$ (the local sum of exponents) and a vector $A_{h}$ (the local attention) per token.
The above formulations provide a simplified conceptual overview. In practice, for efficient inference, we use Flash Attention (Dao, 2024) to attend to the KV cache on each host and apply the log-sum-exp trick from online softmax (Milakov & Gimelshein, 2018) to ensure numerical stability during global attention aggregation.
@@
3 Experiments
We empirically evaluate Star Attention using several Llama-based models across multiple long-context benchmarks with sequence lengths ranging from 16K to 1M tokens, assessing both its accuracy and inference speedup relative to established baselines. We also investigate the accuracy-speed trade-offs as a function of block size and provide a granular breakdown of Star Attention’s effectiveness across different domains. Our results demonstrate that Star Attention consistently achieves near-parity with full global attention in accuracy while delivering substantial speedups, especially on large models and long-context tasks.
@@ 3 Experiments
16 4 92.22 -0.94% 1.1x
32 8 87.53 +1.17% 1.2x
64 16 84.79 -1.42% 1.8x
Llama-3.1-8B-Instruct Meta-AI (2024) 128 32 76.31 -1.90% 2.7x
16 4 95.09 -2.71% 1.7x
32 8 94.61 -2.55% 2.0x
Llama-3.1-70B-Instruct Meta-AI (2024) 64 16 88.54 -1.44% 4.7x
Table 1: Accuracy and relative inference speedup of Star Attention compared to Ring Attention on RULER across sequence lengths from 16K to 128K. Accuracy is reported as the absolute difference from Ring Attention and speedup reflects relative improvements in inference efficiency. Star Attention significantly accelerates inference with minimal accuracy loss.
Figure 4: Accuracy comparison of Star Attention and Global Attention on RULER and BABILong from 16K to 128K sequence lengths using various models. All runs use a block and anchor block size set to one-quarter of the total sequence length. Star Attention maintains 97-100% of the accuracy of global attention, and in some cases, even outperform it.
@@ 3.1 Setup
Models. We conduct experiments using both the base and instruct variants of Llama-3.1 8B which support context lengths up to 128K tokens (Meta-AI, 2024). To evaluate scalability beyond this range, we use gradientai-Llama-3-8B-Instruct-262K and gradientai-Llama-3-8B-Instruct-1048K that extend Llama-3-8B’s context to 256K and 1M tokens respectively (Gradient.ai, 2024). We further assess the impact of model scale using Llama-3.1-70B-Instruct. Across all configurations, Star Attention demonstrates increasing speedup benefits with larger models and longer sequences.
Baseline. We compare Star Attention against three strong baselines: (i) Ring Attention (Liu et al., 2024a), a distributed attention mechanism that computes global block-wise attention by circulating each host’s KV cache in a ring pattern across all the hosts; (ii) StreamingLLM (Xiao et al., 2024b), a sparse attention method that combines global sink tokens with sliding window attention. We use a configuration having 1000 global sink tokens along with a sliding window of 8000 tokens; and (iii) MInference (Jiang et al., 2024), which utilizes three distinct sparse attention patterns, dynamically selecting the optimal pattern per head in an offline search setting. Among these, only Ring Attention is a distributed algorithm designed to scale inference across multiple GPUs. Since Star Attention also targets distributed efficiency, we report speedup metrics relative to Ring Attention, while accuracy comparisons are provided for all three baselines.
Configuration. We implement Star Attention in both HuggingFace Transformers library (Wolf et al., 2020) and NVIDIA’s TRT-LLM framework (NVIDIA, 2023). All experiments are conducted on NVIDIA A100 GPUs with bfloat16 precision. Optimization techniques such as Flash Attention are applied uniformly across Star and Ring Attention implementations to ensure a fair comparison. Reported results are based on the HuggingFace implementation, with similar relative trends observed across TRT-LLM. Additional details regarding our experimental setup can be found in Appendix B.
Evaluation Benchmarks. We evaluate our method on three benchmarks, each testing unique aspects of long context understanding: (i) RULER (Hsieh et al., 2024): a synthetic benchmark with 13 tasks categorized into 4 domains: Needle-in-a-Haystack (Retrieval), Multi-Hop Tracing, Aggregation, and Question Answering. (ii) BABILong (Kuratov et al., 2024): a benchmark of 5 tasks requiring reasoning over multiple supporting facts encoded in the context to generate accurate answers. (iii) InfiniteBench (Zhang et al., 2024): a diverse collection of 10 real-world and synthetic tasks spanning summarization, multilingual QA, code debugging, and retrieval. Further details on the benchmarks and specific tasks can be found in Appendix C.
@@ 3.2 Results
Table 1 presents the accuracy and relative speedup of Star Attention compared to Ring Attention (representing full global attention in a distributed setting) on RULER, across sequence lengths from 16K to 128K tokens. In each setting, the context and the anchor block size are set to one-quarter of the total sequence length. Star Attention maintains high accuracy, typically within 0-3% of global attention, while delivering significant speedups, ranging from 1.1$\times$ to 4.7$\times$, depending on the model size and sequence length. The speedup becomes more pronounced for larger models. For instance, the Llama-3.1-70B-Instruct model exhibits a 4.7$\times$ acceleration at 64K tokens with minimal accuracy drop. This highlights Star Attention’s suitability for high-throughput inference and its ability to preserve model’s accuracy even with a significantly reduced context window.
(a) Accuracy vs. Context Block Size
(b) Accuracy vs. Anchor Block Size
Figure 5: Impact of context and anchor block sizes on the accuracy of Star Attention at 128K sequence length with Llama-3.1-8B Instruct. (a) Accuracy as a function of context block size, with anchor block size matched to it. (b) Accuracy as a function of anchor block size, with context block size fixed at 32K. Larger block sizes yield consistent accuracy improvements, highlighting the benefit of broader receptive fields for long-context understanding.
To evaluate generalization beyond RULER, we benchmark Star Attention on BABILong using Llama-3.1-8B-Base, gradientai-Llama-3-8B-Instruct-262K, and gradientai-Llama-3-8B-Instruct-1048K. As shown in Figure 4, Star Attention consistently achieves near-parity with full attention across all tasks up to 128K sequence length, with an accuracy drop typically below 3%. However, we observe anomalies for the Llama-3.1-8B base model on BABILong, likely due to format-specific generation requirements that challenge non-instruction-tuned models, particularly at longer sequence lengths.
3.3 Comparison with Other Sparse Attention Methods
While our primary comparison focuses on Ring Attention (Liu et al., 2024a) due to its distributed design, we also evaluate Star Attention against two strong non-distributed sparse attention baselines: StreamingLLM (Xiao et al., 2024b) and MInference (Jiang et al., 2024). These methods represent alternative strategies for long-context efficiency under constrained compute budgets and provide complementary perspectives on accuracy trade-offs.
Table 2 reports accuracy on RULER using Llama-3.1-8B-Instruct across sequence lengths from 16K to 128K tokens. Star Attention outperforms both the methods, with the performance gap widening at longer context lengths. Notably, Star Attention maintains accuracy closest to the baseline (full attention) across all settings, demonstrating its robustness in extended-context reasoning.
To assess generalization beyond synthetic tasks, we further evaluate all methods on InfiniteBench. As shown in Table 3, Star Attention achieves the highest average accuracy across 10 diverse tasks spanning summarization, multilingual QA, code debugging, and retrieval. It excels especially in retrieval-heavy tasks such as PassKey, NumRetr, and KVRetr while also delivering competitive results across other categories. These findings highlight Star Attention’s ability to generalize beyond synthetic benchmarks and handle real-world, instruction-heavy tasks with long-range dependencies.
Methods 16K 32K 64K 128K Average
Full Attn. 92.22 87.53 84.79 76.31 85.21
StreamingLLM 74.76 48.56 26.2 30.77 45.07
MInference 93.27 86.54 84.86 58.17 80.71
Star Attention 91.27 88.70 83.37 74.41 84.44
Table 2: Accuracy comparison of different methods on RULER from 16K to 128K sequence length using Llama-3.1-8B-Instruct. Star Attention performs closest to Full Attention and outperforms others at longer sequences.
Methods En. En. En. En. Zh. Code. Math. Retr. Retr. Retr. Avg.
Sum QA MC Dia QA Debug Find PassKey Num KV
Full Attn. 31.91 25.92 69.43 21.5 31.95 16.75 24.29 99.15 99.66 60 48.06
StreamingLLM 30.15 10.15 41.05 8.5 22.38 8.63 17.71 2.71 5.93 0 14.72
MInference 1.0 31.04 22 63.76 14.5 28.7 5.33 27.43 56.78 77.12 14 34.07
Star Attention 31.85 25.92 69 22 30.37 24.37 26.29 93.22 96.27 45.8 46.51
Table 3: Accuracy comparison of different methods on InfiniteBench using Llama-3.1-8B-Instruct. Star Attention performs closest to Full Attention and outperforms others across all the diverse tasks.